OpenStudy (anonymous):

Add or subtract. then simplify by collecting like radical terms, if possible. Assume that all expressions under radicals represent nonnegative numbers. √3-2√108+5√27

OpenStudy (anonymous):

\[\sqrt{3-2} \sqrt{108+5} \sqrt{27}\] This is what it looks like.

OpenStudy (jit4won):

Are you sure ?

OpenStudy (jit4won):

Wanna get it done the way it is ?

OpenStudy (anonymous):

\[\sqrt{3}-2 \sqrt{108}+5 \sqrt{27}\] Sorry i meant like this

OpenStudy (jit4won):

Okay ..Gonna solve

OpenStudy (jit4won):